OpenStudy (queenbb):

What is the slope-intercept form of the linear equation x + 4y = 12?

KjSaif (kjsaif):

first of all what do you think? :)

OpenStudy (waterbird):

1. x + 4y = 12 2. 4y = -x + 12 3. y = -x/4 + 3

KjSaif (kjsaif):

^^

RhondaSommer (rhondasommer):

yep!

OpenStudy (queenbb):

so hats the answer

OpenStudy (queenbb):

whats

OpenStudy (waterbird):

I just wrote out the steps. The third step is the answer.